Anil owns a plot worth Rs. 10,000. He sells it to Mukesh


at a profit of 10o/o. After sometime, Mukesh sells it .back


to Anil at a loss of 10%, then Anil:


(A) loses Rs. 100 (B) loses Rs. 900


(C) gains Rs. 100 (D) gains Rs. 1100

Grade:Upto college level

1 Answers

Abhishek Doyal
10 years ago
ans is (A) loses rs. 100
he sells it to mukesh at 10% profit =10000+10000*10/100=11000
so he sells it to mukesh at=11000
Mukesh sells it back to Anil at a loss of 10% =11000-11000*10/100=9900
so 10000-9900=100
Abhishek doyal